OpenStudy (anonymous):

Need help on verifying the identity! (cosx/1+sinx) + (1+sinx/cosx) = 2secx?

OpenStudy (anonymous):

\[\frac{\cos(x)}{1+\sin(x)}+\frac{1+\sin(x)}{\cos(x)}=2\sec(x)\] \[\frac{\cos(x)\times [\cos(x)]}{[1+\sin(x)]\times [\cos(x)]}+\frac{[1+\sin(x)]\times [1+\sin(x)]}{\cos(x)\times [1+\sin(x)]}=2\sec(x) \]

OpenStudy (anonymous):

can you firstly simplify this?

OpenStudy (anonymous):

good job, you did for him:)

OpenStudy (anonymous):

What did you do to get the first step, @Princer_Jones

OpenStudy (anonymous):

a/b+c/d=(ad+bc)bd this formula @DDCommander

OpenStudy (anonymous):

(ad+bc)/bd

OpenStudy (anonymous):

Sorry for being nitpicky, but what is the formula's name?

OpenStudy (anonymous):

@DDCommander check the attachment

OpenStudy (anonymous):

Thanks a lot, lifesaver